OpenStudy (anonymous):

Find the sum to 5 terms of the geometric progression whose first term is 54 and fourth term is 2.

OpenStudy (tkhunny):

Geometric progression a, a*r, a*r^2, a*r^3, a*r^4, ... We are given a = 54, thus, 54, 54*r, 54*r^2, 54*r^3, 54*r^4, ... We are given 2 = 54*r^3 Where does that leave us?

OpenStudy (anonymous):

ohh i got it :) thanx stupid me